OpenStudy (anonymous):

How can we divide 2.4 by 16 ... Without using calculator ?

OpenStudy (anonymous):

\[\frac{ 2.4 }{ 16 }=\frac{ 24 }{ 160 }=\frac{ 6·4 }{ 40·4 }=\frac{ 6 }{ 40 }=\frac{ 3 }{ 20 }=\frac{ 15 }{ 100 }=0.15\]